# Mixed Precision Iterative Refinement for Linear Inverse Problems

### v1

@@ 4.2 Mixed precision IR on the Tikhonov problem
Before extending Theorem 4.3, we present the generalization of Proposition 1 to the mixed precision case. We provide in Algorithm 2 the steps with their associated precisions for computing preconditioned Landweber. In Proposition 2 we highlight the parts of the filtered solution that are computed in Pr1, Pr2, and Pr3. The framework for the computations of the filter factors is summarized in Algorithm 3.
Algorithm 2 Mixed precision preconditioned Landweber
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x^{(0)}=0\in\mathbb{R}^{n},\,\,\alpha>0$
3  Output: $x^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $M^{T}M=\left(A^{T}A+\alpha^{2}I\right)$ $[\text{in }\text{Pr}_{1}]$
5  for $k=0,1,2,\dots$ do
6        $s^{(k)}=A^{T}\left(b-Ax^{(k)}\right)$$[\text{in }\text{Pr}_{3}]$
7        Solve $(M^{T}M)h^{(k)}=s^{(k)}$ $[\text{in }\text{Pr}_{2}]$
8        $x^{(k+1)}=x^{(k)}+h^{(k)}$ $[\text{in }\text{Pr}_{2}]\,\,$
9  end for
Proposition 2.
(Mixed precision preconditioned Landweber with filter factors)
The $k^{th}$ solution, $x^{(k)}_{PL}$, of (12) with preconditioner $M=D_{M}V_{M}^{T}$ as defined in Section 2 in precisions Pr1, Pr2, and Pr3 may be written as
with filter factors per $k^{th}$ step given by
under the assumption that $V_{M}^{T}V_{A}=I=V_{A}^{T}V_{M}$.
Proof 4.5.
From (12) we have that
$x^{(k)}_{PL}=x^{(k-1)}_{PL}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}.$
Using the definition of $M$, $\hat{A}$, and the SVD of $A$ we can rewrite preconditioned Landweber using our orthogonality assumption as follows
Here, $d_{M,j}^{2}=(\sigma^{2}_{M,j}+\alpha^{2})$ represents the components from the preconditioner computed in Pr1 and whose component-wise action is computed in Pr2. Simplifying, we have
where the update to $\psi_{j}^{(k-1)}$ is also computed in Pr2.
Algorithm 3 Mixed precision filter factors of preconditioned Landweber
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\Sigma^{2}_{M}\in\mathbb{R}^{n\times n},\,\,\alpha>0$
3  Output: $\Psi^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}$ $[\text{in }\text{Pr}_{1}]$
5  for $k=0,1,2,\dots$ do
6        for $j=1,2,\dots,n$ do
9              $\psi_{j}^{(k+1)}=\psi_{j}^{(k)}+g_{j}^{(k)}$ $[\text{in }\text{Pr}_{2}]\,\,$
10        end for
11  end for
With Proposition 2 we now consider the reformulation of IR on the Tikhonov problem in mixed precision so as to determine its corresponding filter factors. We provide the mixed precision variant of Algorithm 1 in Algorithm 4 for reference. Recall that the $k^{th}$ solution of (18) can be written as a recursive relation involving the $k^{th}$ iterate of preconditoned Landweber and the previous $k-1$ iterates of IR on the Tikhonov problem. In Lemma 4.6 we reformulate (18) to write the $k^{th}$ iterate of the IR problem as a filtered solution whose constituent parts are computed in precisions that mimic their partnered forms in Algorithm 4.
@@ 1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $A\in\mathbb{R}^{m\times n},\,\,x^{(0)}\in\mathbb{R}^{n}$
3  Output: $x^{(k+1)}\in\mathbb{R}^{n}$
@@ 4  Compute $M^{T}M=\left(A^{T}A+\alpha^{2}I\right);$ $[\text{in }\text{Pr}_{1}]$
6        $r^{(k)}=b-Ax^{(k)};$ $[\text{in }\text{Pr}_{3}]$
7        $s^{(k)}=A^{T}r^{(k)}-\alpha^{2}x^{(k)};$ $[\text{in }\text{Pr}_{3}]$
@@ 9        $x^{(k+1)}=x^{(k)}+h^{(k)};$ $[\text{in }\text{Pr}_{2}]\,\,$
The $k^{th}$ solution of IR on the Tikhonov problem, $x_{IR}^{(k)}$, in precisions Pr1, Pr2, and Pr3 may be written with respect to the $k^{th}$ solution of mixed precision preconditioned Landweber computed by Proposition 2. Precisely,
(24) $\begin{split}&x^{(k)}_{IR}=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}\left[\left(M^{T}M\right)\left[x^{(k)}_{PL}-x^{(k-1)}_{PL}\right]+\Gamma-\alpha^{2}x_{IR}^{(k-1)}\right]\end{split}$
@@
We now describe the computation of the filter factors of IR on the Tikhonov problem in precisions Pr1, Pr2, and Pr3. The result is provided without proof as the derivations follow in a similar manner to Theorem 4.3 when starting with the result of Lemma 4.1. Here, we adhere to the same convention outlined in Proposition 2 regarding the computation of residuals, linear systems solves, and the preconditioner in appropriate precisions. Algorithm 5 summarizes the framework for computing the filter factors.
@@
with the $j^{th}$ filter factor per $k^{th}$ iterative step given by
$\begin{split}\phi_{j}^{(k)}=\phi_{j}^{(k-1)}+\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}&\Biggl{[}\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\left[\psi_{j}^{(k)}-\psi_{j}^{(k-1)}\right]-\alpha^{2}\phi^{(k-1)}_{j}\\ &+\left(\frac{\alpha^{2}\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-2}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-2-i)}\Biggr{]}\end{split}$
@@ 1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $\Phi^{(0)}=0\in\mathbb{R}^{n},\,\,\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\Sigma^{2}_{M}\in\mathbb{R}^{n\times n},\,\,\alpha>0$
@@ 3  Output: $\Phi^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}$ $[\text{in }\text{Pr}_{1}]$
@@
6        Compute $\Psi^{(k+1)}$ using Algorithm 3 $[\text{in }\text{Pr}_{1},\text{Pr}_{2},\text{Pr}_{3}]$
@@ 5.1 Effective filter factors
To assess the experimental veracity of our filter factor results from Section 4 for IR on the Tikhonov problem we compare against experimentally derived effective filter factors. Using (5), the effective filter factors $\omega^{(k)}_{j}$ of an approximate solution at the $k^{th}$ iteration may be computed as follows
@@ 5.2 Low precision simulation in MATLAB
To ascertain the effectiveness of mixed precision IR on the Tikhonov problem and the experimental validity of the results presented in Section 4 we utilize the software package chop introduced by Higham and Pranesh in [17]. The chop function simulates lower precision arithmetic by rounding array entries given in a MATLAB native precision (e.g. single or double) to a target precision which is stored in a higher precision with non-utilized representation bits set to zero. Chen et al. in [6] considered the application of chop for the solution of structured inverse problems as well as associated numerical considerations of utilizing chop.
@@
Table 1: Precisions considered in the numerical results: shorthand given as an integer, bits given for the mantissa and exponent, and whether precision is IEEE standard or not.
Pr Shorthand Name Exponent Bits Mantissa Bits IEEE Std.
$1$ fp64 $11$ $52$ Yes
$2$ fp32 $8$ $23$ Yes
$3$ fp16 $5$ $10$ Yes
$4$ fp8 $4$ $3$ No (custom)
@@ 5.3 Examples and Results
We compare the solution quality of MP-IR to what we will term the Approximated Iterative Refinement (AIR) method inspired by the work of Donatelli and Hanke in [8]. Here, we let the preconditioner be $M^{T}M=\left(C^{\ast}C+\alpha^{2}I\right)$ where $C$ represents a circulant approximation to $A$. From this structure, the preconditioner is normal and admits a spectral decomposition that can be computed efficiently. We approximate $C\approx A$ using the scheme devised by Chan in [5] which minimizes the Frobenius norm of the difference between $A$ and $C$ amongst a family of circulants. This approximation scheme was analyzed by Strela and Tyrtyshnikov [29] and found to be a good choice from the viewpoint of eigenvalue clustering when the vector to be reconstructed takes on many zero values.
@@
(Spectra - filter factors in fp64) - We begin our filter factor analysis by evaluating the filtering behavior of MP-IR on the Spectra problem contaminated by $1\%$ noise in fp64. For this experiment we chose the regularization parameter $\alpha^{2}=1e-2$; in the restoration part of this example we expand on this choice. Figure 1 displays the filter factors from Theorem 4.3, Theorem 4.8, and the effective filter factors compared to those of Tikhonov when considering all computations in fp64 after $1$ iteration. We comment that there is no visual discernible difference amongst the plots thereby experimentally verifying that: (1) the filter factors of Theorem 4.3 and Theorem 4.8 coincide when all floating point computations are done in fp64, i.e. when $(\text{Pr}_{1},\,\text{Pr}_{2},\,\text{Pr}_{3})=(1,1,1)$ and (2) that all three sets of filter factors match those of Tikhonov. Table 2 (discussed below) numerically confirms the agreement between Theorem 4.8 and the effective filter factors in fp64 working precision.
@@
Figure 1: Filter factors of the Spectra problem with $1\%$ noise compared against those of Tikhonov with $\alpha^{2}=1e-2$ after $1$ iteration using (a) the effective filter factors, (b) Theorem 4.3, and (c) Theorem 4.8.
[…]
@@
Figure 2: MP-IR filter factors (bottom row) determined by Theorem 4.8 with $(\text{Pr}_{1},\text{Pr}_{2},\text{Pr}_{3})=(3,2,1)$ for $1\%$ noise compared to their entry wise computations of $|\Phi^{(k)}-\Omega^{(k)}|$ (top row) at iteration numbers (a) 1, (b) 5, and (c) 10.

### v2

@@ 4.2 Mixed precision IR on the Tikhonov problem
We now consider the reformulation of IR on the Tikhonov problem in mixed precision so as to determine its corresponding filter factors. In three precisions the $k^{th}$ solution of (19) may formulated as
(25) $x^{(k)}_{IR}=\underbrace{x^{(k-1)}_{IR}+{\underbrace{\vphantom{\Big{|}}\left(M^{T}M\right)}_{\text{Pr}_{1}}}^{-1}\underbrace{\left[A^{T}r^{(k-1)}-\alpha^{2}x^{(k-1)}_{IR}\right]}_{\text{Pr}_{3}}}_{\text{Pr}_{2}}$
where we utilize braces to visually denote the usage of mixed precision. This mixed precision variant of Algorithm 1 is provided in Algorithm 2 for reference. Recall that the $k^{th}$ solution of (19) can be written as a recursive relation involving the $k^{th}$ iterate of preconditoned Landweber and the previous $k-1$ iterates of IR on the Tikhonov problem. In Lemma 4.5 we reformulate (19) to write the $k^{th}$ iterate of the IR problem as a filtered solution whose constituent parts are computed in precisions that mimic their partnered forms of (25) and Algorithm 2.
@@ 1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $A\in\mathbb{R}^{m\times n},\,\,x_{IR}^{(0)}\in\mathbb{R}^{n}$
3  Output: $x_{IR}^{(k+1)}\in\mathbb{R}^{n}$
@@ 4  Compute $M^{T}M=\left(A^{T}A+\alpha^{2}I\right);$ $[\text{in }\text{Pr}_{1}]$
6        $r^{(k)}=b-Ax_{IR}^{(k)};$ $[\text{in }\text{Pr}_{3}]$
7        $s^{(k)}=A^{T}r^{(k)}-\alpha^{2}x_{IR}^{(k)};$ $[\text{in }\text{Pr}_{3}]$
@@ 9        $x_{IR}^{(k+1)}=x_{IR}^{(k)}+h^{(k)};$ $[\text{in }\text{Pr}_{2}]\,\,$
The $k^{th}$ solution of IR on the Tikhonov problem, $x_{IR}^{(k)}$, in precisions Pr1, Pr2, and Pr3 may be written with respect to the $k^{th}$ solution of preconditioned Landweber, $x^{(k)}_{PL}$. Precisely,
(26) $\begin{split}&x^{(k)}_{IR}=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=\underbrace{x^{(k-1)}_{IR}+{\underbrace{\vphantom{\Big{|}}\left(M^{T}M\right)}_{\text{Pr}_{1}}}^{-1}\underbrace{\left[\left(M^{T}M\right)\left[x^{(k)}_{PL}-x^{(k-1)}_{PL}\right]+\Gamma-\alpha^{2}x_{IR}^{(k-1)}\right]}_{\text{Pr}_{3}}}_{\text{Pr}_{2}}\end{split}$
@@
We comment that during the computation of the $k^{th}$ iterate of (26) that the product $\left(M^{T}M\right)\left[x^{(k)}_{PL}-x^{(k-1)}_{PL}\right]$ is computed in Pr3 but $\left(M^{T}M\right)$ is computed in Pr1. Thus, before presenting the mixed precision generalization of Theorem 4.3, we present in Proposition 2 the variable precision case of Proposition 1. In Algorithm 3 we provide the steps for computing preconditioned Landweber where we note that the preconditioner $\left(M^{T}M\right)$ is an input and the working precision is Pr3 (as per Lemma 4.5). The framework for the computations of the filter factors is summarized in Algorithm 4.
Algorithm 3 Variable precision preconditioned Landweber
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x_{PL}^{(0)}=0\in\mathbb{R}^{n},\,\,\alpha>0$,   $M^{T}M$ (in Pr1)
3  Output: $x_{PL}^{(k+1)}\in\mathbb{R}^{n}$
4  for $k=0,1,2,\dots$ do
5        $s^{(k)}=A^{T}\left(b-Ax_{PL}^{(k)}\right)$$[\text{in }\text{Pr}_{3}]$
6        Solve $(M^{T}M)h^{(k)}=s^{(k)}$ $[\text{in }\text{Pr}_{3}]$
7        $x_{PL}^{(k+1)}=x_{PL}^{(k)}+h^{(k)}$ $[\text{in }\text{Pr}_{3}]\,\,$
8  end for
Proposition 2.
(Variable precision preconditioned Landweber with filter factors)
The $k^{th}$ solution, $x^{(k)}_{PL}$, of (13) computed in Pr3 with preconditioner $M=D_{M}V_{M}^{T}$ as defined in Section 2 such that $M^{T}M$ is computed in Pr1 may be written as
with filter factors per $k^{th}$ step given by
under the assumption that $V_{M}^{T}V_{A}=I=V_{A}^{T}V_{M}$.
Proof 4.7.
From (13) we have that
$x^{(k)}_{PL}=x^{(k-1)}_{PL}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}.$
With the definition of $M$, $\hat{A}$, and the SVD of $A$ we can rewrite preconditioned Landweber using our orthogonality assumption as follows
Here, $d_{M,j}^{2}=(\sigma^{2}_{M,j}+\alpha^{2})$ represents the components from the preconditioner computed in Pr1 and whose component-wise action is computed in Pr3. Simplifying, we have
where the update to $\psi_{j}^{(k-1)}$ is also computed in Pr3.
Algorithm 4 Filter factors of variable precision preconditioned Landweber
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\alpha>0$,
3              $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\in\mathbb{R},\quad\text{for }j=1,2,\dots,n$ (in Pr1)
4  Output: $\Psi^{(k+1)}\in\mathbb{R}^{n}$
5  for $k=0,1,2,\dots$ do
6        for $j=1,2,\dots,n$ do
8              $g_{j}^{(k)}=\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}q_{j}^{(k)}$ $[\text{in }\text{Pr}_{3}]$
9              $\psi_{j}^{(k+1)}=\psi_{j}^{(k)}+g_{j}^{(k)}$ $[\text{in }\text{Pr}_{3}]\,\,$
10        end for
11  end for
We comment that the filter factors $\psi^{(k)}_{j}$ for $j=1,2,\dots,n$ computed by Proposition 2 are equivalent to those of Proposition 1 in exact arithmetic. This can be observed by using relations (16) and (29) to consider the difference between subsequent filter factor iterates, i.e., $\psi^{(k)}_{j}-\psi^{(k-1)}_{j}$. It’s worthwhile to mention that that the recursive nature of the filter factors from Proposition 2 is a product of the structure that arises from Algorithm 3.
We now describe the computation of the filter factors of IR on the Tikhonov problem in precisions Pr1, Pr2, and Pr3. The result is provided without proof as the derivations follow in a similar manner to Theorem 4.3 when starting with the result of Lemma 4.1. Here, we adhere to the same appropriate precision choices described at the start of the section in (25) with respect to the computation of residuals, linear systems solves, and the preconditioner. Algorithm 5 summarizes the framework for computing the filter factors.
@@ 9              $\psi_{j}^{(k+1)}=\psi_{j}^{(k)}+g_{j}^{(k)}$ $[\text{in }\text{Pr}_{3}]\,\,$
with the $j^{th}$ filter factor per $k^{th}$ iterative step with indicated precision given by
(30) ${\color[rgb]{0,0,0}\phi_{j}^{(k)}=\underbrace{\phi_{j}^{(k-1)}+{\underbrace{\vphantom{\Big{|}}\left(\sigma_{M,j}^{2}+\alpha^{2}\right)}_{\text{Pr}_{1}}}^{-1}\underbrace{\Biggl{[}\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\left[\psi_{j}^{(k)}-\psi_{j}^{(k-1)}\right]+\Delta-\alpha^{2}\phi^{(k-1)}_{j}\Biggr{]}}_{\text{Pr}_{3}}}_{\text{Pr}_{2}}}$
where
$\Delta=\left(\frac{\alpha^{2}\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-2}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-2-i)}$
@@ 1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $\Phi^{(0)}=0\in\mathbb{R}^{n},\,\,\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\alpha>0$
@@ 3  Output: $\Phi^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\in\mathbb{R},\,j=1,\dots,n$ from $M^{T}M=\left(A^{T}A+\alpha^{2}I\right)$ $[\text{in }\text{Pr}_{1}]$
@@
6        Compute $\Psi^{(k+1)}$ using Algorithm 4 $[\text{in }\text{Pr}_{3}]$
@@ 5.1 Effective filter factors
To assess the experimental veracity of our filter factor results from Section 4 for IR on the Tikhonov problem we compare against experimentally derived filter factors which we term effective filter factors. Using (5), the effective filter factors $\omega^{(k)}_{j}$ of an approximate solution at the $k^{th}$ iteration may be computed as follows
@@ 5.2 Low precision simulation in MATLAB
To ascertain the effectiveness of mixed precision IR on the Tikhonov problem and the experimental validity of the results presented in Section 4 we utilize the software package chop introduced by Higham and Pranesh in [18]. The chop function simulates lower precision arithmetic by rounding array entries given in a MATLAB native precision (e.g. single or double) to a target precision which is stored in a higher precision with non-utilized representation bits set to zero. Chen et al. in [6] consider the application of chop for the solution of structured inverse problems and investigate associated numerical considerations of its usage.
@@
Table 1: Precisions considered in the numerical results: shorthand given as an integer, bits given for the mantissa and exponent, the approximate decimal round-off error, and whether precision is IEEE standard or not.
Pr Shorthand Name Exponent Bits Mantissa Bits Approx. Decimal Round-off IEEE Standard
$1$ fp64 $11$ $52$ $-161.11e\text{-}$ 16 Yes
$2$ fp32 $8$ $23$ $-85.96e\text{-}$ 8 Yes
$3$ fp16 $5$ $10$ $-44.88e\text{-}$ 4 Yes
@@ 5.3 Examples and Results
To efficiently use chop in an experimental setting to simulate low precision arithmetic, the matricies $A$ we utilize herein for the numerical examples are Toeplitz structured, which, for inverse problems are often well approximated by circulant matrices; see e.g., [1, 8]. Because of this, we compare the solution quality of MP-IR to what we will term the Approximated Iterative Refinement (AIR) method where we use circulant approximations of $A$ within the preconditioner $M^{T}M$ to construct a normal preconditioning matrix which admits a spectral decomposition that can be computed efficiently. Specifically, we let the preconditioner be $M^{T}M=\left(C^{T}C+\alpha^{2}I\right)$ where $C$ represents a circulant approximation to $A$. We approximate $C\approx A$ using the scheme devised by Chan in [5] which minimizes the Frobenius norm of the difference between $A$ and $C$ amongst a family of circulants. This approximation scheme was analyzed by Strela and Tyrtyshnikov [30] and found to be a good choice from the viewpoint of eigenvalue clustering when the vector to be reconstructed takes on many zero values.
@@
(Spectra - filter factors in fp64) - We begin with a comparison of our filter factor results presented herein by evaluating the filtering behavior of MP-IR on the Spectra problem contaminated by $1\%$ noise in fp64. For this experiment we chose the regularization parameter $\alpha^{2}=1e\text{-}2$; in the restoration part of this example we expand on this choice. Figure 1 displays the filter factors from Theorem 4.3, Theorem 4.8, and the effective filter factors compared to those of Tikhonov (24) when considering all computations in fp64 after $1$ iteration. We comment that there is no visual discernible difference amongst the plots thereby experimentally verifying that: (1) the filter factors of Theorem 4.3 and Theorem 4.8 coincide when all floating point computations are done in fp64, i.e. when $(\text{Pr}_{1},\,\text{Pr}_{2},\,\text{Pr}_{3})=(1,1,1)$ and (2) that all three sets of filter factors match those of Tikhonov. Table 2 (below) numerically confirms the agreement to unit round-off between Theorem 4.8 and the effective filter factors in fp64 working precision.
@@
Figure 1: Spectra example: filter factors of standard Tikhonov (24) compared against filter factors of iterative refinement computed using: (a) the effective filter factors (31), (b) the filter factors computed by Theorem 4.3 (23), and (c) the filter factors computed by Theorem 4.8 (30). The filter factors are compared using $\alpha^{2}=1e\text{-}2$ after $1$ iteration with $1\%$ noise.
(Spectra - filter factors in mixed precision) - We utilize Theorem 4.8 to investigate the filter factors of MP-IR and compare them against the effective filter factors that come from the iterates computed by Algorithm 2 where again we utilize a noise level of $1\%$ and $\alpha^{2}=1e\text{-}2$. In Table 2 we provide the summary statistics of the absolute difference between $\Phi^{(k)}$ and $\Omega^{(k)}$ of the 64 entries computed by (31). We note that for most precision combinations investigated, that by the $5^{th}$ iteration the mean error had converged to the unit round-off of the precision associated with Pr1 (see Table 1). We illustrate in Figure 2 that compared to the behavior summarized in Figure 1 of filter factors computed in fp64 that if the precision combination is changed to $(\text{Pr}_{1},\text{Pr}_{2},\text{Pr}_{3})=(3,2,1)$ that the variation of the mean entry-wise value of $|\Phi^{(k)}-\Omega^{(k)}|$ takes on approximately the unit round-off of the precision associated with the preconditioner (i.e., Pr1). This can be observed in the variation of sequential entries in the top row of Figure 2. Furthermore, we observe in the bottom row of the figure that the absolute largest deviation between the effective filter factors and those computed by Theorem 4.8 arise in the first iteration for index numbers corresponding to a rapidly changing derivative. We surmise these differences may be a result of the orthogonality assumption made throughout or of the choice of the regularization parameter.
@@
Figure 2: MP-IR filter factors (bottom row) determined by Theorem 4.8 with $(\text{Pr}_{1},\text{Pr}_{2},\text{Pr}_{3})=(3,2,1)$ for $1\%$ noise compared to their entry-wise computations of $|\Phi^{(k)}-\Omega^{(k)}|$ (top row) at iteration numbers (a) 1, (b) 5, and (c) 10.